# Your Paper Title Here

### v1

@@ 2.1 Linear Calibration Correction (LCC)
For LCC, we approximate the relationship between EO-ML predictions and ground-truth wealth on the held-out calibration set by a linear transformation:
$\hat{Y}_{i}=kY_{i}+m,$
@@
where $0<k\leq 1$ reflects shrinkage toward the mean and $m\in\mathbb{R}$ is the intercept, representing the mean fixed bias. This linearity might not always be empirically satisfied, but occurs regularly in our data (see, e.g., Figure A.III.3). When these predictions are used in place of true outcomes to estimate causal effects, this shrinkage attenuates estimated treatment effects.
@@
Let $\tau=\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]$ denote the true Average Treatment Effect (ATE). Using $\hat{Y}_{i}$ instead of $Y_{i}$ for ATE estimation yields the bias shown in Proposition 2.1.
Proposition 2.1 (Attenuation bias under linear shrinkage)
Let $Y_{i}$ be the true outcome and $\hat{Y}_{i}$ an ML prediction obeying the linear calibration model:
$\hat{Y}_{i}\;=\;kY_{i}+m+e_{i}\quad(0<k\leq 1,\;m\in\mathbb{R}),$
@@
with $k$ and $m$ constants, and $e_{i}$ mean 0 noise. Let $A_{i}\in\{0,1\}$ denote a binary treatment that is randomly assigned ($A_{i}\perp Y_{i}(a)$ for $a\in\mathcal{A}$). Define the true average treatment effect (ATE) $\tau:=\mathbb{E}[Y_{i}(1)]-\mathbb{E}[Y_{i}(0)]$ and the estimate, $\hat{\tau}$, from difference in predicted expected outcome:
$\hat{\tau}=\mathbb{E}[\hat{Y}_{i}\mid A_{i}=1]-\mathbb{E}[\hat{Y}_{i}\mid A_{i}=0].$
@@
Proof sketch.
Take expectations using linearity:
@@
$=\mathbb{E}[kY_{i}+m\mid A_{i}=1]-\mathbb{E}[kY_{i}+m\mid A_{i}=0]=k\tau.$
Replacing observed with potential outcomes follows from randomization. The estimated ATE is biased downward by a factor of $k$.
Now, to adjust for this bias, we will estimate $k$ and $m$ on a held-out calibration set from the training data by regressing $\hat{Y}_{i}$ on $Y_{i}$ (i.e., fitting the linear model $\hat{Y}_{i}=\hat{k}Y_{i}+\hat{m}+e_{i}$, where $e_{i}$ is the residual). We then invert this fitted relationship (solving for $Y_{i}$) to obtain a calibrated estimate, given by $\hat{Y}^{L}_{i}=\frac{\hat{Y}_{i}-\hat{m}}{\hat{k}}$, assuming residuals $e_{i}$ with mean zero conditional on covariates, independent of treatment $A_{i}$. This calibrated estimator restores the scale of the outcome variable, yielding Proposition 2.2:
Proposition 2.2 (Consistency of the LCC-corrected estimator)
Under the conditions of Proposition 2.1, estimate $(k,m)$ on an independent calibration sample and set:
@@
[…]
Proof sketch.
[…]
@@
$=(k/\hat{k})Y_{i}+(m-\hat{m})/\hat{k}+e_{i}/\hat{k}\to Y_{i}+e_{i}/k$
Hence,
$\hat{\tau}^{L}\to\mathbb{E}[Y_{i}+e_{i}/k\mid A_{i}=1]+\mathbb{E}[Y_{i}+\epsilon_{i}/k\mid A_{i}=0]=\tau,$
@@
since $\mathbb{E}[e_{i}]=0$ and under under noise independence, $e_{i}\perp A_{i}$. A drawback is that we assume a linear model for the shrinkage; estimating $\hat{m}$ and $\hat{k}$ introduces another potential source of error.
2.2 Tweedie’s Correction for Debiasing
To go beyond the linearity assumption in LCC, we now turn to Tweedie’s formula, which allows us to correct shrinkage bias in machine learning predictions by estimating the true outcome as the prediction plus a term involving the derivative of the log-density of the predictions. We follow Efron (2011), who revised the classical results in the context of selection bias as an automatic bias-correction tool in an empirical Bayes framework. We leverage the same core insight here to adjust for regression-to-the-mean bias in satellite-imputed outcomes when analyzing causal effects.
Setup. We here treat the shrinkage mechanism probabilistically:
[…]
$\varepsilon_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2}),\;\;\varepsilon_{i}\;\perp\;(Y_{i},A_{i}).$
The factor $k<1$ pulls extreme $Y_{i}$ values toward the center, lowering $\operatorname*{\mathrm{Var}}(\hat{Y}_{i})$ whenever $\sigma^{2}$ is not too large. Setting $k=1$ recovers the additive-noise model sometimes used in empirical-Bayes texts (Efron, 2012), so (1) nests the classic case.
This noise model enables us to leverage Tweedie’s formula to recover the posterior mean of the latent outcome, accounting for shrinkage without requiring linearity assumptions.
Posterior mean via a scaled Tweedie identity.
Write $W_{i}:=kY_{i}$. Then $\hat{Y}_{i}=W_{i}+\varepsilon_{i}$ with the same $\varepsilon_{i}$ as above, and the Tweedie score identity (Efron 2011) [^1] For any $Z=\theta+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, $\operatorname*{\mathbb{E}}[\theta\mid Z=z]=z+\sigma^{2}\,\partial_{x}\!\log p_{Z}(z)\,$. gives
$\operatorname*{\mathbb{E}}\!\bigl{[}\,W_{i}\mid\hat{Y}_{i}=\hat{y}\,\bigr{]}\;=\;\hat{y}\;+\;\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}}(\hat{y}),$ (2)
where $p_{\hat{Y}}$ is the marginal density of $\hat{Y}_{i}$; Dividing by $k$ recovers the latent outcome:
${\;\operatorname*{\mathbb{E}}\!\bigl{[}\,Y_{i}\mid\hat{Y}_{i}=\hat{y}\,\bigr{]}\;=\;\frac{\hat{y}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}}(\hat{y})}{k}.\;}$ (3)
For causal estimation, the relevant posterior is conditional on treatment: $\operatorname*{\mathbb{E}}[Y_{i}\mid\hat{Y}_{i}=\hat{y},A_{i}=a]=\frac{\hat{y}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}=a}(\hat{y})}{k}$, since the noise is independent of $A_{i}$ but treatment shifts the distribution of $Y_{i}$.
By expressing the posterior mean of the scaled latent variable in terms of the score function, we obtain a convenient identity that underlies the Tweedie adjustment.
Debiased point estimate.
Equation 3 motivates the Tweedie-adjusted outcome:
$\widetilde{Y}_{i}\;:=\;\frac{\hat{Y}_{i}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}}(\hat{Y}_{i})}{k},$ (4)
which depends solely on the prediction $\hat{Y}_{i}$, treatment $A_{i}$, and two nuisance parameters, $k$ and $\sigma^{2}$.
With this adjustment defined, we now establish its unbiasedness properties under the assumed model.
Proposition 2.3 (Unbiasedness)
Under model (1), $\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}\mid\hat{Y}_{i},A_{i}]=\operatorname*{\mathbb{E}}[Y_{i}\mid\hat{Y}_{i},A_{i}]$ and hence by iterated expectations, $\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}]=\operatorname*{\mathbb{E}}[Y_{i}]$.
Proof Sketch.
Take conditional expectations in (4) and apply the group-specific version of (3) and iterated expectations.
Having established the unbiasedness of the Tweedie-adjusted outcome both conditionally and unconditionally, we now leverage this property to demonstrate that a simple difference-in-means estimator based on these adjusted outcomes yields an unbiased estimate of the average treatment effect under random assignment.
Proposition 2.4 (Tweedie-corrected ATE)
Let $\hat{\tau}^{T}:=\hat{\operatorname*{\mathbb{E}}}[\widetilde{Y}_{i}\mid A_{i}=1]-\hat{\operatorname*{\mathbb{E}}}[\widetilde{Y}_{i}\mid A_{i}=0],$ where $\hat{\operatorname*{\mathbb{E}}}$ denotes the sample mean. Under (1) and random assignment, $\operatorname*{\mathbb{E}}[\hat{\tau}^{T}]=\tau$.
Proof Sketch.
$\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}\mid A_{i}=a]=\operatorname*{\mathbb{E}}[Y_{i}\mid A_{i}=a]$ by Prop. 2.3; subtract.
Figure 3: Illustration of Tweedie’s correction mechanism, showing how the posterior mean adjusts for shrinkage bias in predicted outcomes under Gaussian noise.
Estimation
The parameters $k$ and $\sigma^{2}$ come from a held-out calibration set $\mathcal{I}_{\text{cal}}\subset\mathcal{I}_{\text{Train}}$ upstream; scores are estimated downstream:
  1. 1.
    Shrinkage factor $k$. Regress $\hat{Y}_{i}$ on $Y_{i}$: $\hat{Y}_{i}=\hat{k}\,Y_{i}+\hat{m}+\text{resid}$. We keep the slope $\hat{k}$ (discarding the intercept, absorbed into $Y_{i}$).
  2. 2.
    Noise variance $\sigma^{2}$. Use the residuals: $\hat{\sigma}^{2}=|\mathcal{I}_{\text{cal}}|^{-1}\sum_{i\in\mathcal{I}_{\text{cal}}}(\hat{Y}_{i}-\hat{k}Y_{i})^{2}.$
  3. 3.
    Group-specific score functions $\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}=a}(\hat{y})$. Downstream, for each $a\in\{0,1\}$, fit a smooth density $\widehat{p}_{\hat{Y}\mid A_{i}=a}$ to $\{\hat{Y}_{i}:i\in\mathcal{I}_{\text{Trial}},A_{i}=a\}$. Trim or regularize the derivative where $\widehat{p}_{\hat{Y}\mid A_{i}=a}$ is small.
The operational formula passed downstream is in the end:
$\widehat{\widetilde{Y}}_{i}=\frac{\hat{Y}_{i}+\hat{\sigma}^{2}\,\partial_{\hat{y}}\log\widehat{p}_{\hat{Y}\mid A_{i}}(\hat{Y}_{i})}{\hat{k}}.$
Relation to LCC.
When the score term tends to zero (e.g., occurring when the variance on outcomes, $Y_{i}$, grows large), $\widetilde{Y}_{i}=\hat{Y}_{i}/k$—exactly an LCC correction. Tweedie’s adjustment, therefore, refines LCC with a local, data-driven addend that vanishes in regions where $p_{\hat{Y}\mid A_{i}}$ is flat and grows where shrinkage is most severe. For simplicity, we will, for all empirical Tweedie analyses henceforth, set $k=1$ to isolate the performance contribution of the local, data-driven adjustment (versus that from the LCC-type correction).
3 Experiments with Simulated Data

### v2

@@ 1. Linear Calibration Correction (LCC)
For LCC, we approximate the relationship between EO-ML predictions and ground-truth wealth on a held-out calibration set by a linear transformation:
$\mathbb{E}[\hat{Y}_{i}\mid Y_{i},\,A_{i}]=\mathbb{E}[\hat{Y}_{i}\mid Y_{i}]=k\,Y_{i}+m,\;\;k>0,\ m\in\mathbb{R}.\;(A1)$
@@
When $0<k\leq 1$, shrinkage toward the mean is indicated; $m\in\mathbb{R}$ is the intercept, representing the mean fixed bias. The first equality essentially indicates that there is now conditional independence of $\hat{Y}_{i}$ from group/treatment variable $A_{i}\in\{0,1\}$ given $Y_{i}$; the second equality indicates a global affine (linear) calibration of predictions to truth, i.e., $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}]=kY_{i}+m$ with constant slope and intercept. This linearity might not always be empirically satisfied, but appears to occur approximately in our data (see, e.g., Figure A.III.3). When these predictions are used in place of true outcomes to estimate causal effects, this shrinkage attenuates estimated treatment effects.
@@
Let $\tau=\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]$ denote the true Average Treatment Effect (ATE). Using $\hat{Y}_{i}$ instead of $Y_{i}$ for ATE estimation yields the bias shown in Proposition 1.
Proposition 1 (Attenuation bias under linear shrinkage)
@@
Let $Y_{i}$ be the true outcome and $\hat{Y}_{i}$ an ML prediction obeying the linear calibration model (A1) with $k$ and $m$ constants. Let $A_{i}\in\{0,1\}$ denote a binary treatment that is randomly assigned ($A_{i}\perp Y_{i}(a)$ for $a\in\mathcal{A}$). Define the true average treatment effect (ATE) $\tau:=\mathbb{E}[Y_{i}(1)]-\mathbb{E}[Y_{i}(0)]$ and the estimate, $\hat{\tau}$, from the difference in predicted expected outcomes:
$\hat{\tau}:=\mathbb{E}[\hat{Y}_{i}\mid A_{i}=1]-\mathbb{E}[\hat{Y}_{i}\mid A_{i}=0].$
@@
Proof. Take expectations and use linearity:
@@
$=\mathbb{E}[\mathbb{E}[\hat{Y}_{i}\mid Y_{i},A_{i}=1]\mid A_{i}=1]$
$\qquad-\mathbb{E}[\mathbb{E}[\hat{Y}_{i}\mid Y_{i},A_{i}=0]\mid A_{i}=0]$
$=\mathbb{E}[kY_{i}+m\mid A_{i}=1]-\mathbb{E}[kY_{i}+m\mid A_{i}=0]$
$=k\{\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]\}=k\tau.$
Replacing observed with potential outcomes follows from randomization; ATE estimates with $\hat{Y}_{i}$ are biased downward by a factor of $k$.
Now, to adjust for this attenuation bias, we will estimate $k$ and $m$ on a held-out calibration set from the training data by regressing $\hat{Y}_{i}$ on $Y_{i}$ (i.e., fitting the linear model with functional form, $\hat{Y}_{i}=\hat{k}Y_{i}+\hat{m}+e_{i}$, where $e_{i}$ are mean 0 residuals). We then invert this fitted relationship (solving for $Y_{i}$, which gives $Y_{i}=\frac{(\hat{Y}_{i}-\hat{m}-e_{i})}{\hat{k}}$) to obtain a calibrated estimate, given by $\hat{Y}^{L}_{i}=\frac{\hat{Y}_{i}-\hat{m}}{\hat{k}}$, assuming conditional mean zero residuals $e_{i}$. Here, $\hat{Y}^{L}_{i}$ reflects the calibrated estimate of the true outcome, $Y_{i}$. This calibrated estimator restores the scale of the outcome variable, yielding Proposition 2:
Proposition 2 (Consistency of the LCC‑corrected estimator)
[…]
@@
If $(\hat{k},\hat{m})\xrightarrow{p}(k,m)$, and $k>0$, then $\hat{\tau}^{L}\xrightarrow{p}\tau$; i.e., the linear calibration correction removes the attenuation bias.
Proof sketch. By Slutsky’s theorem, since $(\hat{k},\hat{m})\xrightarrow{p}(k,m)$, it follows that:
@@
$=\underbrace{(k/\hat{k})Y_{i}}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\to Y_{i}}+\underbrace{(m-\hat{m})/\hat{k}}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\to 0}+\underbrace{e_{i}/\hat{k}}_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}e_{i}/k}\to Y_{i}+e_{i}/k.$
Since $\mathbb{E}[e_{i}|A_{i}]=0$ under the conditional mean 0 assumption outlined earlier,
$\hat{\tau}^{L}$ $\to\mathbb{E}[Y_{i}+e_{i}/k\mid A_{i}=1]-\mathbb{E}[Y_{i}+e_{i}/k\mid A_{i}=0]$
$=\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]=\tau.$
@@
A drawback is that we assume a linear model for the shrinkage mechanism, which enters through an assumed functional form for the expected value of $\hat{Y}_{i}$ given $Y_{i}$ (cf. A1); estimating $\hat{m}$ and $\hat{k}$ introduces another potential source of error.
2. Tweedie’s correction
To go beyond the linearity assumption in LCC in a way that respects the structure of ML predictions, we now turn to Tweedie’s identity, which provides a local, density–score–based way to reverse shrinkage without assuming global linearity of the measurement error process.
A first instinct might be to formalize shrinkage in a way that generalizes beyond the linearity assumption of LCC by using a measurement-error model, in which the prediction is a noisy version of the truth:
$\hat{Y}_{i}=Y_{i}+\varepsilon_{i},\;\;\;\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$
However, in ML predictions scenarios, models approximate $\mathbb{E}[Y_{i}\mid\mathbf{X}_{i}]$ for covariates $\mathbf{X}_{i}$ and thus predictions tend to exhibit systematically lower variance than targets due to shrinkage toward the mean (the opposite of what is implied in the baseline classical case, where $\textrm{Var}(\hat{Y}_{i})>\textrm{Var}(Y_{i})$). Indeed, upstream ML models are trained to minimize expected loss over $\mathbf{X}_{i}$ (shalev2014understanding):
$Y_{i}=\hat{Y}_{i}+\epsilon_{i},\quad\hat{Y}_{i}=f_{\theta}(\mathbf{X}_{i})$
where $\theta$ are neural parameters optimized to reduce MSE. Any residual noise $Y_{i}-\hat{Y}_{i}$ thus arises after the prediction, approximately independent of $\hat{Y}_{i}$ itself in distribution (as the model would otherwise exploit correlations to reduce loss). This reasoning therefore inverts the classical view, motivating adoption of the Berkson error model (carroll2006measurement; heid2004two):
$Y_{i}=\hat{Y}_{i}+\varepsilon_{i},\;\;\;\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})\perp\hat{Y}_{i},$
as a more natural framing for debiasing ML surrogates.
Adapting Tweedie’s formula.
To reverse shrinkage under this model without fresh labels, we propose an approach using Tweedie’s formula, introduced by robbins1956empirical and commonly encountered in Gaussian diffusion models:
  • Tweedie’s Formula:
  • For any random variable, $Z$, with $Z=\omega+\varepsilon$,
    where $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$,
  • $\mathbb{E}[\omega\mid Z=z]=z+\sigma^{2}\,\frac{d}{dz}\log p_{Z}(z)$   (efron2011tweedie).
In its native form, Tweedie’s formula under the Berkson model gives us an expression for $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}=y]$, i.e., recovering model predictions under attenuation bias from true labels. Our aim is to derive a novel debiasing method that works in the opposite direction, i.e., recovering estimates of the true labels without attenuation bias in downstream applications. We achieve this goal by introducing a new-label-free corrected pseudo-outcome, $\widetilde{Y}_{i}$, whose conditional mean recovers the truth, $\mathbb{E}[\widetilde{Y}_{i}\mid Y_{i}]=Y_{i}$. To obtain this end, we need to deploy the score-swapping relation:
$\frac{d}{dy}\log p_{Y}(y)=\mathbb{E}\left[\frac{d}{d\hat{y}}\log p_{\hat{Y}}(\hat{Y}_{i})\mid Y_{i}=y\right],$
as derived in Appendix I, to the Tweedie identity:
$\mathbb{E}\left[\hat{Y}_{i}\mid Y_{i}=y\right]$ $=y+\sigma^{2}\frac{d}{dy}\log p_{Y}(y)$
$=y+\sigma^{2}\,\mathbb{E}\left[\frac{d}{d\hat{y}}\log p_{\hat{Y}}(\hat{Y}_{i})\mid Y_{i}=y\right]$
$\Longrightarrow y$ $=\mathbb{E}\left[\hat{Y}_{i}-\sigma^{2}\frac{d}{d\hat{y}}\log p_{\hat{Y}}(\hat{Y}_{i})\mid Y_{i}=y\right],$
where the move from the second to the third line follows from linearity of expectations.
Intuitively, subtraction in the correction arises from inverting the forward application of Tweedie’s formula under the Berkson model. Unconditionally, $\mathbb{E}[\hat{Y}_{i}]=\mathbb{E}[Y_{i}]$, so predictions are centered around the true values on average (as ensured by MSE minimization during training). Conditionally, however, $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}=y]=y+\sigma^{2}\frac{d}{dy}\log p_{Y}(y)$, where the score term attenuates extreme predictions toward the population mean—for instance, when $y$ exceeds the mode of $p_{Y}$, the score is negative and pulls the conditional expectation below $y$ and towards the mode. In other words, while the overall distribution of predictions matches that of the truths in expectation, a given prediction (absent adjustment) shrinks toward the mean conditional on the realized true value. The inversion thus requires subtracting the score—now estimated via the swapped relation on $p_{\hat{Y}}$—to reverse this attenuation and expand the tails.
This discussion then yields the Tweedie corrected pseudo-outcome under the Berkson model:
$\widetilde{Y}_{i}=\hat{Y}_{i}-\sigma^{2}\,\frac{d}{d\hat{y}}\log p_{\hat{Y}}(\hat{Y}_{i}).$ (1)
The score term expands both tails: when the score is positive (where we evaluate the score rising to the left of a local mode), it pushes $\hat{Y}_{i}$ downward; when the score is negative (score evaluated to the right of a local mode) it pushes $\hat{Y}_{i}$ upward; around a mode, where the score is near zero, adjustment is negligible. In practice, $\sigma$ and the score function will need to be estimated on data: $\sigma$ is estimated from residuals; the score function with, e.g., a KDE on $\{\hat{Y}_{i}\}_{i=1}^{n}$.
Proposition 3 (Tweedie calibration identity)
The pseudo-outcome satisfies: $\mathbb{E}\!\left[\widetilde{Y}_{i}\,\middle|\,Y_{i}\right]\;=\;Y_{i}.$
Proof sketch. By Tweedie’s identity, $\mathbb{E}[\hat{Y}_{i}\mid Y_{i}=y]=y+\sigma^{2}\,\frac{d}{dy}\log p_{Y}(y)$. By score‑swapping (Appendix I), $\frac{d}{dy}\log p_{Y}(y)=\mathbb{E}[\frac{d}{d\hat{Y}}\log p_{\hat{Y}}(\hat{Y}_{i})\mid Y_{i}=y]$. Take conditional expectations of $\widetilde{Y}_{i}$ given $Y_{i}=y$ to obtain $y$.
Remark. This is the key statement giving rise to well-calibrated downstream applications of the “one-map” ML prediction. Note that the variance of $\tilde{Y}_{i}$ will almost certainly be larger than the variance of $\hat{Y}_{i}$, so for purely predictive purposes, the overall MSE of pseudo-outcomes may also be larger that of $\hat{Y}_{i}$; however, downstream teams will for causal inference purposes be concerned with well-calibrated group conditional means—discussed later.
Proposition 4 (Local Reduction to LCC)
If $p_{\hat{Y}}$ is locally Gaussian with mean $\mu$ and variance $\gamma^{2}$, then the score is $(\hat{y}-\mu)/\gamma^{2}$ and the pseudo‑outcome is the affine map:
$\widetilde{Y}_{i}$ $\;=\;\hat{Y}_{i}-\sigma^{2}\left(-\frac{\hat{Y}_{i}-\mu}{\gamma^{2}}\right)=\mu+\Bigl(1+\tfrac{\sigma^{2}}{\gamma^{2}}\Bigr)(\hat{Y}_{i}-\mu),$
i.e., a linear calibration with slope $(1+\sigma^{2}/\gamma^{2})$ increasing in the ratio of measurement error variance to local signal variance that expands extremes away from $\mu$.
Proof. Plug the Gaussian score into $\widetilde{Y}_{i}$.
Further Relations to LCC.
When the score term is negligible (e.g., flat density), $\widetilde{Y}_{i}\approx\hat{Y}_{i}$, reverting to uncorrected predictions. While LCC provides a single global scaling adjustment for errors that operates in a uniform manner, Tweedie adjustment here can handle more complex distributional patterns, rescaling depending on the position of predictions relative to local (possibly multiple) modes. Thus, LCC captures a single regression-to-the-mean dynamic, while Tweedie adjustment captures a broader range of error processes, adjusting for regression to local modes.
Proposition 5 (Unbiasedness of Tweedie means)
Under the assumptions of Proposition 3, for any partitioning variable $A_{i}$ such that $\hat{Y}_{i}\perp A_{i}\;|\;Y_{i}$ (which implies $\tilde{Y}_{i}\perp A_{i}\;|\;Y_{i}$ [because $\tilde{Y}_{i}$ is a deterministic function of $\hat{Y}_{i}$]):
$\mathbb{E}[\widetilde{Y}_{i}\mid A_{i}=a]\;=\;\mathbb{E}[Y_{i}\mid A_{i}=a].$
Consequently, the difference‑in‑means estimator $\hat{\tau}:=\hat{\mathbb{E}}[\widetilde{Y}_{i}\mid A_{i}=1]-\hat{\mathbb{E}}[\widetilde{Y}_{i}\mid A_{i}=0]$ satisfies $\mathbb{E}[\hat{\tau}]=\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]$. Under random assignment, $\mathbb{E}[\hat{\tau}]=\tau$.
Proof sketch. Iterated expectations using Proposition 3 yield $\mathbb{E}[\widetilde{Y}_{i}\mid A_{i}=a]=\mathbb{E}\{\mathbb{E}[\widetilde{Y}_{i}\mid Y_{i},A_{i}=a]\mid A_{i}=a\}=\mathbb{E}[Y_{i}\mid A_{i}=a]$. Independence of $A_{i}$ and $\epsilon_{i}$ is key here, since it implies that there is no additional relationship between $\tilde{Y}_{i}$ and $Y_{i}$ once $Y_{i}$ is known, so $\mathbb{E}[\tilde{Y}_{i}|Y_{i}]=\mathbb{E}[\tilde{Y}_{i}|Y_{i},A]$. Randomization identifies $\mathbb{E}[Y_{i}\mid A_{i}=a]=\mathbb{E}[Y_{i}(a)]$ and thus enables unbiased ATE estimates in RCT scenarios.
Remark. The key assumption here, $\hat{Y}_{i}\perp A_{i}\;|\;Y_{i}$, indicates that, conditional on the true value of $Y_{i}$, there is no additional signal relating $A_{i}$ and $\hat{Y}_{i}$. For example, if $A_{i}$ is gender and $Y_{i}$ is income, the assumption states that, among truly high-income individuals, the model error does not predict gender. This unbiasedness result does therefore not mean all forms of bias are eliminated, simply that the bias added from attenuation is eliminated. See Appendix I for a discussion of plug-in consistency, which is, in practice, complicated by how overfitting in the training set can lead to distribution shifts affecting performance.
Viewed through an optimization lens, the Tweedie pseudo-outcome $\widetilde{Y}_{i}=\hat{Y}_{i}-\sigma^{2}\frac{d}{d\hat{Y}}\log p_{\hat{Y}}(\hat{Y}_{i})$ is exactly a single gradient-descent step on the log-density $\log p_{\hat{Y}}$, with learning rate $\eta=\sigma^{2}$, that nudges predictions away from the over-shrunk high-density region toward the tails (see Appendix I). This perspective naturally points to a broader class of Tweedie-style updates obtained by varying the step size, number of steps, and implicit error model. In the Appendix, we discuss how one might organize the resulting Tweedie corrections as a family indexed by the implicit error model that links the latent outcome $Y_{i}$ and the ML prediction $\hat{Y}_{i}$. The variants differ by the structure one is willing to assume and by how one wishes to correct global scaling (via a single calibration slope $k$ and number of update steps) in addition to local shrinkage. In our experiments below, we focus on the Berkson variant we have been analyzing—leaving to future work elucidation of the full range of bias-variance trade-offs in the Tweedie family.
Experiments with Simulated Data